DFR Telecoms Diary
1 January :
3 January :
4 January : I got the keys and charger from Norchard and went to the Junction signal box to see how the exchange was. The battery volts were down to 33. I did the usual, I felt all the uniselectors and sure enough an incoming selector was stuck and quite warm. I think an intermittent fault on the cable had been driving the uniselector off normal and home continually until the volts reached the point where the uniselector stuck on. I disconnected both incoming junctions and homed the uniselector manually. The volts promptly went up into the forties and was still climbing when I left.
6 January : Peter, Roger, Sam and I met at Parkend mostly for a look at progress on the work there and a bit of a general discussion on future jobs and procedures. It looks as though we will get the cable for the Town Crossing job next week. Peter, Roger met Jean and I for lunch at Kaplan's.
10 January : Sam, Rick, Martin and I were going to have a go at finding the DP3Y to DP3X cable fault. Both 301C testers refused to switch on. Perhaps they hate the cold. In the end we got the 301B going and went to Whitecroft to try that. Useless. We could hardly make out the cursor on the CRT screen, it's too small and if the sun shines it makes it extremely difficult. I walked the length but could not find an obvious disturbance. The only thing that might be a pointer was a bit of a signal pointing to 70 metres away where there is sound whistle sign. Could it have been moved? Sam will enquire. So we have ended up really no wiser after a day tramping the lineside. The 301Cs are now in my home waiting for me to check inside but are no better after a night in the warm. Peter and Charles carried on with the Parkend job. Martin and Rick had a bit of a go at some maintenance.
13 January : I went to the railway to see if the town crossing cable was going to be laid. The answer - no! It's still not been ordered. I had a chat with Bob Hawker trying to persuade him to find us a decent cabinet for parkend DP1E. He will have a go and he also hopes to find concrete posts at Middle Forge where S and T are to replace a cabinet with another box.
17 January : Sam and I went to the Whitecroft area to continue looking for the cable fault. Rick and Martin stayed in the exchange to test with us. After a lot of wandering up and down the line, the 301C leads us to believe that the fault is in the section DP3Y to DP3Z and about 12 m from DP3Y. We either have to dig and find the fault for confirmation or assume the 301C is right and dig further along to find the cable, insert a cabinet and lay a new cable over the faulty length. We shall be glad to see the back of this cable fault. The token is now not working so the cable is all faulty except possibly for one pair. Sam put the token out of use. Rick and Martin did some maintenance in Norchard and Peter and Charles continued work on the Parkend rewire.
19 January : Sam and Peter agreed to meet in the cafe at Norchard at 11:30, Peter had an early lunch and after meeting with Sam who had to attend a meeting arranged to meet Sam at Parkend at 2:30 PM. Peter proceeded to Parkend with a view to removing DP 3 cable from the old MDF and reconnecting it to the new MDF by way of one of the newly provided Box Conn 254's. Recovery of cable proceeded well until it came to removing it from behind the bench / chest of drawers in the exchange, it would not budge and was stuck fast. Peter cleared all of the items currently on the bench with the thought of how to move the chest of drawers beneath to release the cable stuck behind. After laying on the floor the problem was found to be that the cable retaining clips were acting as a lock every time the cable was pulled. Fortunately when the cable clips were fixed some one had used cross point screws, at arms length with a 150mm cross point drill bit Peter was able to remove the offending cable clips and the cable was recovered, coiled up so as to retain spare length and terminated. Literately within minutes of Peter finishing Sam arrived and it was decided this was a good point to stop for the day.
20 January : Peter, Sam and Roger worked at Parkend today. The goal was to sort out the cable between the Exchange and the Station and to start extend the broadband internet connection from the station building to the exchange and signal box. With the aid of 3 different drills (including the 110v hammer drill) and several visits to the loft by Roger and Sam, we had the BT master socket moved, the external bell on the BT line re-cabled, the 10 pair from the exchange re-routed via the loft and some old cabling stripped out.
24 January : As the day was continually rainy, there was no chance of working on our cable network problems. Consequently we all had quite a good and productive day. Rick and Martin cleared a long standing fault on a final at Norchard. In the end it was silver migration between U9 and 11 preventing the selector from releasing. They are fitting another tubular heater to that rack as all the migration faults have been on that equipment. Peter, Sam and Charles worked on the Parkend recabling job. It's approaching completion. I made a start on simplifying the Parkend ringer control arrangements but could not finish phase 1 as I was unable to obtain access to the back of the rack due to the large amount of gear hanging around in the exchange. First job will be to start on clearance of the stores in the exchange. Roger P says that the town crossing cable will be here in a fortnight.
26 January : Peter and Sam surveyed the state of telecoms along the line and upcoming projects. This included the new MDF at Lydney Junction, as well as having a quick look at Whitecroft (which needs documenting properly).
27 January : Peter and Sam ran in most of the cables into the signal box, with the concentrator rack-head cable, cables to telephones along the back wall and cables to the S&T racking. A little more work is needed for this before they are all terminated. Steve Harris then constructed some "Heritage" wooden trunking around where they rise out of the floor in the signal box. Peter and Sam then went to Caerphilly to collect an eBay purchase described as 130 sip phones, as a job lot (which turned out to be 170!), before depositing them into the telecoms office to await sorting. A quick sift through revealed several types, and a number of them were sorted into a box, which was delivered to Roger for testing/evaluation.
28 January : Peter and Sam met about 09.30 to start sifting through the new SIP telephones, cleaning and testing them before they were boxed up ready for storage. Peter left to do other things, and Sam was joined by Jane Bambridge, who had come in to help. By the end of the day, all 170 phones were cleaned, boxed and two thirds of them were put away in the attic of the shop office.
31 January : Peter, Sam, Rick, Charles and Martin met in the telecoms office, and we were joined by Pete Hyneman, ex-BT, who was interested in joining the team. Rick and Martin returned to their maintenance jobs in the Norchard exchange, while Sam, Peter, Charles and Pete went to Parkend to work on various tasks. Sam showed Pete around the site, to look at what wiring needed completing. On arrival, Peter joined us in the exchange, while Charles went to complete the earthing in DP3.
4 February : Sam, Peter, Roger, myself and Alan met in the cafe. Alan is a possible new recruit, ex BT but radio trained rather than telephones. He joined us for the morning to get an idea of our work. We all went off to Parkend. Sam and Roger spent the day sorting out the programming side of the Asterisk with the SIP phones we have obtained. Peter did several small jobs at Parkend towards the rewiring job. I completed the work changing over to the new ringer control relays. I installed the new clock/NU Spare Levels relay set and got it wired in and tested. It worked. I also recovered the relay strip of the old ringer control though the wiring has to come out later. We had lunch at Kaplan's.
7 February : Sam and Charles reorganised the E side pairs at Parkend DP1 with Peter in the exchange. I walked the line to check on the plastic bag DP1E, it seems OK for now. I also got three temporary floodlights fitted to enable us to work on the Parkend North exchange. Ray had a go at a puzzling fault on a Parkend linefinder. He has got so far in determining where the intermittent fault is but it still needs further investigation. Rick and Martin changed the ringer in Norchard. The spare works OK but doesn't sound as good as the regular ringer. This now needs new brushes, which we should have somewhere.
14 February : Rainy, cold and a chill wind. We met in the cafe as we can hardly get into our office at the moment with stores all over the place. Ray and Martin went off to spend their day at Parkend trying to locate a fault on a first selector. In the end it seems to be silver migration again. This will be the first instance of this at Parkend. The selector is now working well but we will keep an eye on it.
17 February : Sam, Roger and I worked on terminating the replacement cable between DP3Y and DP3Z. We got everything terminated. Sam and I had a disagreement about what constitutes pairs in the railway colour coded cable we used, and so we have ended up with split pairs in one or other of the cables each side of the new DP3YA. A first check of the circuits between Norchard and Parkend showed them all to be disconnected somewhere. Sam and Roger wandered off up the line having an initial look to see if things had gone wrong at the damaged DP1E or if the recent permanent way work at Parkend has caused further problems. I await their report.
21 February :
Rick and Martin spent the day on strowger maintenance. Peter, Ray and our new man, Alan, spent the day at Parkend.
Sam, Charles and I walked the route between the "tin can joint" near DP3F to the Town crossing to inspect the cabinets and see progress on the trenching. It's coming along nice and deep.
24 February :
Peter and Sam started terminating cables on the Parkend North Krone frame. They terminated all of the operating floor cables, and the cables to the signalling rack.
28 February : Peter M and Charles worked on Parkend North, working on terminations on the frame, and a few small jobs. Rick and Martin worked in the exchange at Norchard, continuing their maintenance overhaul. Sam and Pete H went to look at the faults between Whitecroft and Norchard, testing DP3Y, 3X and 3V. In the afternoon, they reterminated several joints in DP3S, hoping that would clear the fault, which it did not.
7 March : Peter M and Charles returned to Parkend to work on the new MDF. Rick, Ray and Martin repaired the Signal Post Telephone at Middle Forge, this turned out to be a simple case of the screw terminal block having come undone. They then went around testing and checking all of the other SPTs. They found signal 13 to have a faulty ring. Sam and Allen, with John's testing expertise helping in the Exchange, went to investigate the faults between Parkend and Whitecroft. This again started in DP3S, where we discovered that the cable between DP3S and DP3A had been terminated wrongly when installed some time ago! We were then able to reconnect a Norchard outgoing junction, so that calls can now be made across it to Parkend!
10 March : At Parkend Exchange Peter and Roger spent the day installing CAT5 cabling and terminating it in the data cabinet. This gave us 4 network sockets on the bench for testing SIP Phones and ATAs. We installed a PoE switch and rack shelf and powered up the data extenders ready for the Parkend Asterisk server. Rick was lineside testing cabling from Parkend to Whitecroft.
14 March : We met as normal in the cafe. Tea was noticeably shorter than usual, and people were keen to get out and working (wow!). Rick and Martin went fault hunting on the Parkend-Norchard Junctions, with Peter M in the exchange. Sam, Allen, Pete H and Charles went south to sort out the newly laid cable. Pete and Charles went to Lydney town and started terminating the south end of the cable, while Sam and Allen went to Goatfield to work on the north end. By lunchtime, we almost had both ends of the new cable stripped and ready to be terminated in the new cabinet. After lunch, Rick and Martin went off fault tracing again, and then onwards to Parkend with Peter M. Peter and Allen went to work on terminating the new cable in the new cabinet, while Sam and Charles went to find a suitable piece of the old 50pr to make into a temporary tie cable to the new one. By the end of the day we had both the new and the tie pair terminated, with only a few minor issues, which Pete H was kind enough to come and fix the next day. Rick, Martin and Peter managed to get a junction in each direction to Parkend working.
17 March : Sam and I decided to test our junctions, particularly to Lydney Signal Box. This required us to visit the box where we got the two outgoing junctions to Norchard working. We unplugged the incoming circuits but neither would work. We returned to Norchard to find strange conditions coming apparently from the two circuit distributor relay set. In the end I bypassed the set for the two outgoing circuits to Lydney Junction to find that there were still problems. Both junctions had been reversed in the cable work and were locking up to battery on the "wrong" leg. One junction was also "dis" on the earth leg and this fault has to remain for a later date. Rick came in to help after calling him for information. Reversing the good junction got it working. We ended with three out of four junctions back in service. We also have three out of four junctions working to Parkend, so that the only current loss of service would be for calls from Parkend to Lydney Junction. The distibutor circuit now needs to be carefully examined (perhaps for silver migration) on a non running day. Sam went to the town crossing to reconnect the phones there but could not get both to work. Again something odd for Wednesday next.
21 March : Rick and Martin went off chasing the faulty Parkend junctions. I stayed in Norchard exchange testing with them. They finished the day with all four circuits back in service. The alarm circuits from Parkend and Lydney Junction still need to be faulted. Charles went off to Lydney Junction to advise the group replacing the crossing gates about the electrical supplies available. Sam had a day exploring the loft while he, Sam, Alan and Ray got the cables in to the new PA speaker positions. We ended the day with something quite positively completed.
24 March : Peter, Roger and Sam held a meeting on the train about data and Roger showed us some examples of what work he has been doing. This was followed by an early lunch. Peter then went to survey the MDF, while Roger and Sam had a quick look at the raspberry pi asterisk. Roger and Sam then went down to Lydney Junction to look at the drum of fibre there. We found it is loose tubed 32 core fibre, and about 200m. Could be interesting to use it for a test! We returned to Norchard and helped Peter identify all of the cables terminated on the frame. This was followed by some work on the emergency phone.
28 March : I spent the morning at Lydney junction sorting out the alarm/clock circuits. Eventually they worked and I got NU back on 624. Peter and Alan went to parkend to work on the data sockets. After lunch I went to Parkend with Peter to check on what I can do next week when I hope to start on the installation of Parkend North exchange. Rick and Martin chased the Parkend alarm/clock circuit and eventually got it as far as leaving DP3S. Sam and Charles got the Town station phone working again and got stores packed away in the loft.
28 March : Sam, Roger and Peter had a day in the office trying out various data equipment and fibre optic patch leads.
4 April : We had a large group to start the day with tea and cakes in the cafe. Rick and Martin went to do maintenance work on Norchard exchange. Paul and I went to Parkend and started on the job of bringing Parkend North into service. By the end of the day the equipment was functioning and the cables to the MDF tied in and stripped ready for terminating on the exchange rack. Peter, Sam and Pete worked on various jobs around Parkend. Charles went off to be a gatekeeper at the Town crossing for the day. Some of us finished the day at the Parkend Postage Stamp cafe.
7 April : Paul popped in to spend the afternoon with his head in the back of the 10+50, working out how far the restoration project had got, and identifying the next steps. Most of the hard work has already been done, it's mostly tidying left to do now. Paul left a copy of the to-do list in the back of the 10+50 in case someone else gets to it before he has a chance to. Sam and Roger were working on their data project creating a digital PABX in the Telecoms office using Asterisk on a Raspberry Pi computer. The digital PABX has 2 normal phones (via a Grandstream ATA) and 1 digital phone (from Avaya) and can make calls to (and receive from) the Strowger. I was in briefly for a cup of tea and having a look at the caboosh to see what we may need at the next THG swapmeet.
11 April : Sam and I sorted clocks in the caboosh. We may have six that could clean up nicely. We tried for lunch in the cafe but they could not accommodate us as they had a party coming in. We tried the Fountain only to get the same response. We ended up in the Royal Oak who were happy to see us but they had to go and find a chef. Eventually after a too long wait we got a set of very good lunches. I took the clocks home to do them on my bench. I cleaned the best looking two first. They came up nicely but I then found that one of them had an intermittently short circuit coil that needs to be changed. We still have a couple of clocks in the caboosh with broken cases. Hopefully I will get a coil from one of those. Martin and Paul did a bit of strowger maintenance and then went off looking for our "dis" circuits between Norchard and Parkend. Peter and Charles went to Parkend where Charles was looking to provide an extra 13 amp socket for S&T. Sam recovered two wall phones and bell from the original gateside cabinet at Whitecroft, 318 and 597. Peter and Sam, once Peter had refreshed himself in the cafe and Sam had finished serving in the shop, then found the fault preventing the emergency phone from working. They tested that the only level allowed to the phone is level 9.
14 April : Peter and I were at Parkend. Sam joined us for a coffee in the Bike Hire cafe and stayed for a while. Peter was sorting the cables going to the S and T rack and upstairs. I got the cables from the MDF terminated on the exchange rack. Rick looked in to discuss cable records with Peter.
18 April : Rick and Martin set about another attempt to trace the clock and alarm pair. They successfully proved the pair to Whitecroft, and then all the way from there to the Parkend MDF. Peter M assisted with this, alongside updating the labels on the MDF. Allan and Ray rewired 512 to run via DP17. Sam and Pete H connected the newly installed Lydney Town treadle, and tested it. Pete H and Allan had to leave after lunch, so Peter, Rick and Martin went to Parkend to attempt to trace the alarm/clock pair fault from the MDF to the relay set without success as the records of the IDF were not up to date.. Ray and Sam tested that 512 was working, spoke about digital phones and did some wipe-ups.
20 April : It was another visit by the Telecoms Group and their partners to the Forest of Dean student's lunch. This time it was a very substantial Italian meal we enjoyed.
21 April : Nice to see the STARS display at Norchard showing all green for the first time for months. Rick and I sorted out the clock circuit so that the clocks are now working at Parkend but then discovered that two junctions between Norchard and Parkend are not working. Rick checked that it was not more trouble in DP3S but then had to leave to go home, and then away on holiday. The junctions remain busied out. Sam and Pter worked on the cabling in the signal box. We all had lunch at Kaplan's and then went home.
25 April : I worked in Parkend Strowger tidying up. I recovered the fuse mounting 8064s as they have been thrown spare by Peter's work. That made access to the rest of the blocks so much easier. I tried to put the three junctions from Parkend to Parkend North through on the frame but found problems with the records. I will finish this off next time. Peter and Martin worked on the cabling to S and T in the signal box. Sam, Alan and Charles worked on fitting clocks in the Parkend buildings.
28 April : I went to try to jumper the Parkend North junctions in Parkend exchange. I did not get round to it again as we had a contact between two exchange lines. The only place they came together, we thought, was on the new MDF but the fault eventually was traced to the permanent terminations on the tie cable to the strowger. A bit difficult. Peter, Sam and Roger were looking at the signal box requirements. We went to lunch at Kaplan's. Peter, Sam and Roger went to the Junction to survey for fibre optic cable to the station building. I returned to Norchard and set about cleaning up the scrap copper to get it and a load of cable ready for the scrap yard.
2 May : Only Sam, Martin and myself in today. Sam and Martin went off to change out the terminal block in the SPT on the west loop at the Junction. I collected up all the scrap cable and copper and went to meet them. We went to Bendall's to sell the scrap and got an excellent price for it. Copper must be high at present. After lunch Sam and Martin replaced a couple of SPTs.
5 May : I went to Parkend to jumper up the junctions to Parkend North, reinstall the spare incoming selectors and change the exchange phone from 311/390 to 321/390.
9 May : Sam, Peter, Charles, Alan and I spent the day sorting out the Parkend to Whitecroft and Norchard circuits. We ended the day with all of them working with the exception of 333 which appears to be "dis" quite close to Parkend. It's only the lineside test circuit and we may never have connected it through in the past. The two "dis" junctions worked when checked with a tele butt but showed high resistance when tested with a digital meter. We assume that a few volts will not break down any corrosion but the 50 volts, and particularly any back EMFs from the A relays, will hack their way through corrosion. So the faults are gone and cannot be traced at present while the sun shines on the joints. We still have a "dis" junction going south to Lydney Signal Box.
12 May : Roger, Peter and Sam spent the day at Parkend. Peter and Sam connected in the cable from Parkend to Parkend North. Roger worked on the Asterisk digital PBX and set up 2 Avaya digital SIP phones and 2 analogue phones with a Grandstream ATA linked to a local Raspberry Pi running the Asterisk Server software. Calling between phones and calling to and from Norchard was all working, including calling through to Strowger numbers. Then gremlins struck and calls from Parkend Asterisk to Norchard stopped working (but Norchard to Parkend Asterisk calls still worked). It will be looked at next week.
15 May : Sam and Rick went to the Festiniog Railway to pick up some telecoms equipment. They came away with :
19 May : Waiting notes
23 May : We spent the morning mostly on putting away the spares that came from the Festiniog Railway. The caboosh is crammed. After lunch Sam, Rick and I surveyed the cabinets between Upper Forge and Tuft's Bridge for updating and renewal work. We reached the following conclusions :
30 May : It was wet all day so we looked for indoor jobs. Rick and Martin made a start on replacing some PA speakers. Sam and I saw the General Manager to discuss getting electrical work done, making small purchases and our desire to discontinue the 222 fire alarm arrangement. I went to Parkend to make a start on the concentrator cabling to the head. Complicated. Charles was involved in electrical work. Sam and Peter came to Parkend later and got the lines from Parkend North connected and tested. Nice to hear the exchange working. We still have to connect the junctions. Signwriting and labelling is also necessary. We finished the day with coffee at the Royal Oak in Whitecroft.
2 June : Sam, Peter and I met at the Parkend Cycle Cafe for tea. Sam and Peter went surveying cabinets on the Parkend to Whitecroft section and then the cable at the St Mary's footbridge. The footbridge is to be removed shortly to go for refurbishment at Mabey's. I connected up the junctions in Parkend North. The three to Parkend worked straight away. I then went to Norchard to get the records so that I could chase the Norchard level 7 junction down the line. I met Sam and Peter there so I picked up Jean for lunch at Kaplan's. After lunch I went to Whitecroft where I checked the junction through to DP3Z. Whilst there I was told not to park in the entrance and block access to the gates. With two cars in the space I was apparently blocking access for emergency vehicles. Where to park at Whitecroft; inside the gate perhaps? That will be a bind. I went back to Parkend to check that all was well. Sam and Peter turned up. Peter was adding jumpers to the MDF. Sam and I checked the dialling into Parkend North. All the modern pulse phones worked fine, but dial phones give wrong numbers. We did not find the answer but I assume that the extra component in the routing is the auto-auto relay set in Parkend. I will have that out on our next visit and give it a good thrashing. Tea back at the cafe and away quite late at five. A busy day.
6 June : Rick, Martin and I went to Parkend, mostly to file some records and switch on Parkend North. We returned to Norchard to see if the level 7 junction appeared at DP3S. It did not, neither was it in DP3T. We went to Whitecroft to check that it was still arriving there. It was but we decided to switch the routing to a different pair to Norchard in the hope that that would be through. It was not when we returned to Norchard. We went to Upper Forge to DP3U. The circuit was not there so we are dis once more in the Upper Forge to Whitecroft section. Next week for walking through that section. Rick and Ray also had a look at the suspect auto auto relay set at Parkend. The A relay definitely adds pulses. They also found that the relay had a white label coil, so that is the main suspect. I will change the relay for a proper red label A relay. Charles spent the day providing an additional mains socket for the proposed data cabinet in the staff room. Sam and Peter also spent their day at Parkend.
12 June : Rick and Martin came in to reterminate and upgrade DP3U at Upper Forge. The work was done but subsequently we found that some circuits were not now working.
13 June : Rick, Martin and I spent the day sorting out the circuits to Parkend and got everything back to where they should be. We found it so much better with me in the exchange and Rick and Martin at the cabinet. Charles worked on the museum token machines. Peter was not feeling too good and did not get round to any work.
16 June : I went to Parkend and did a bit more on the concentrator wiring. I think it's all connected at the concentrator end. Then I tried the outgoing junctions from Parkend only to find the second outgoing junction was dis out. I put a marking 6 volt battery across the pair to aid testing.
18 June : Rick and I went to the railway to try to sort the dis junction. I went on to Whitecroft to find that I was not getting the 6 volts battery from Parkend. Going on to Parkend I found that I was not getting the battery at DP1, yet it seemed to be leaving the MDF. I met Rick at DP3U where he had replaced the crimps with a connector. I returned to Parkend but we could not see why things were wrong. In the end, after lunch, Rick went back to DP3U to redo the crimps. He went on to Parkend but without a good result. Peter and Sam turned up and a little later so did I. We discussed the situation and I left Sam and Peter to chase the 6 volt battery out to DP1. They got it there in the end but are not sure what cleared the problem, though they did clean the connections as they went. Later Sam rang me at home via the junction. Thank goodness. We do though need a quicker way of testing junctions than trying to find them on the Krone MDF and inserting some wire leads to gain access. We should consider putting the junctions through S & T links in the exchange. That would speed things up considerably, and it would help if we obtained the correct leads for intercepting a circuit on the Krone connectors. We do of course still have DPs to bring up to standard on that route.
19 June : Sam met Pete H at Parkend, with Mike Hardy from the Pway group, who he had found at Norchard looking for something to do. They started by pulling the new cables (20pr and cat5e data cables) through into the station, suffering a couple of setbacks on the way. They resolved these issues, and managed to get them through to the DP, and terminated. They also finished pulling a new 20pr and 2x cat5 from the signal box to the goods shed, this not only provides data, but also allows more circuit growth. They then converted 322 on the platform to a modern wall phone, provided an ATA powered phone in the cabinet alongside it and tested all the phones in the station building as they had disconnected and reconnected them. They all worked. The clock was reset. Finally, they got a SIP phone working over one of the data cables to the signal box.
20 June : Rick, Martin and I had a depressing day chasing faulty pairs between DP3X and DP3Z. How often now have we worked over this part of the line? We found the pairs we were looking for but at the end of the day we had somehow disconnected both outgoing junctions from Parkend and had not cleared the original fault. We have concluded that all our problems have arrived with the use of crimps. We need to reassess our use of the little horrors as we seem to be quite unable to get all our pairs working at any one time. We are also beginning to suspect that the DP3X to DP3Y cable may have been damaged. At least the weather was kind, not too hot, not too wet, and we got to wave at the train frequently. Charles continues with electrical work at Parkend. Sam and Peter continued with the additional cabling work at Parkend.
23 June : Peter and Ian from civils/small works came in today for the afternoon. They had intended to prove some pairs for the new Whitecroft asterisk lines along with clock and signal post phone circuits. This work has now been put off for another day. Sam contacted Peter to ask that a digital line be provided in Whitecroft station. This should have been a simple job of connecting 4310 through on a new piece of temporary cable at the station as the circuit had been proved previously to the cabinet outside the station building. Alas not so, on testing the pair it was dis. The circuit had been disconnected at Parkend in the data cabinet so it should have been just a case of plugging in the right patch cable to the appropriate ATA. No such luck, the ATA configured for 4310 was now configured with 4725. Not having a laptop or passwords etc, the easiest but not right solution was to push 4725 temporarily to Whitecroft. On returning to Whitecroft it was all working with 10 minutes to spare.
25 June : Alan and Sam purchased two fenceposts and erected the first adjacent to but slightly to the Norchard end and closer to the track at DP3X. We've accidentally been given some ten foot posts. Sam plans on leaving a large amount out of the ground and then to see if he can get some high-luminosity paint, to make it more visible!
27 June : Another hot day, getting to 28C. Rick and Martin continued to work on the new PA speaker outside the back of the Severn and Wye Room. I went to Parkend and changed the two A relays in the auto auto relay set for red label types. They seem to be working OK. After lunch we had a disagreement about the best way forward in clearing the troubles on the Norchard to Whitecroft cable. This left Sam to try to clear the faults by recrimping, though we do intend to replace two cabinets, hopefully finding slack cable we can cut off to recrimp on uncorroded wires. Peter and I went back to the Parkend signal box to discuss how to get the cable to the concentrator head. Peter took the head and its shelf home to produce the necessary holes for feeding the cable. Later he e-mailed to say the holes had been done. Charles is continuing to provide additional mains sockets in the Parkend staff room, ready for the data cabinet installation.
28 June : I got called out by Sam to attend an alarm at Lydney Junction. By the time I reached Norchard to pick up the keys, the alarm had gone. It looks as though the mains to the S and T equipment was switched off for a period exceeding the six minute display delay. I went to have a look but all was functioning normally.
29 June : The Telecoms Group attended a Turkish evening at the Jasmine Cafe in Cinderford arranged by Peter. A good event worth running again. We made £95 for the Fund, also a good outcome. Thanks Peter.
4 July : Charles spent the day on electrics at Parkend. Rick and Martin are working their way through the spares recently obtained and are cataloguing them and storing them in Norchard exchange where possible. At the same time the aim is to tidy and make better use of the space in the exchange. I went to Parkend and spent the day reorganising the outlets on the first selector banks to enable a second multi level auto auto relay set to be fitted which will have a digital junction to Norchard. The resultant trunking diagram is shown below.
8 July : Sam and I went to inspect the damge done by Rennes when it set off a lineside fire at Middle Forge yesterday. Our cables are definitely well singed but the armouring has saved them. We seem to have escaped serious damage as the circuits in the cable are testing OK. The Fire Brigade were called to extinguish the blaze.
12 July : Rick and Martin got the SPT working at Middle Forge. They found that the drop wire feeding it had been cut. A clean cut for no obvious reason. Ray did some odd jobs around Norchard. Charles finished off the mains supplies to the proposed data cabinet at Parkend station. Peter and I tried to connect the existing multi level relay set at Parkend to an ATA and Norchard. It would not work. It seems that the pulse ratio of the initial digit is incorrect. After lunch I went home to see what the pulse generating relay in the set I am building would be like. Oh dear, it turns out to be 80% break and 20% make. No good at all. Changing capacitors, altering residuals etc made changes to the pulse speed but not the ratio. That appears to be built into the relay itself. Back to Atkinson's which confirmed this theory. I have therefore now to find a suitable relay. I have a 500 ohm +1500 ohm relay, as per Atkinson, labelled PS. Looks promising.
14 July : I went to Parkend to collect the multi level relay set. I brought it home to modify the pulse generating relay circuit. Worked nicely on the bench. Jean and I went to lunch at Kaplan's with Sam and Peter. They came back to see the relay set chattering away. We took the set back to Parkend and proceeded to test it with a digital junction. We had Roger watching the procedure remotely. The relay set is generating pulses too quickly, probably because I set it up at 50 volts and Parkend mostly runs closer to 55 volts. However, the ATA counts less digits than are dialled, but worryingly this also happens when the regenerator sends out a train of pulses. Yet a butt across the set output is received correctly. I need to think about this. If a modification would help I can include it in the new relay set that I am building. I can, for example, delay the sending of the first digit after seizure a little more in case the ATA is slow on the uptake. I am dialling with a zero ohm loop, so I am unlikely to be able to improve that. Peter and Sam were in trying to list work required to be done. The electrician is nearly finished in the signal box so hopefully we will be able to leave Parkend North powered up and connected to the network apart from the Norchard to Parkend North junction which awaits a cable pair.
18 July : Rick decided to find some pairs in the Norchard to Whitecroft cable. The faults were in the crimps again. He made the repair using the twisting wire and screw connector method. He now has both outgoing junctions from Norchard to Parkend working. Peter and Alan worked on more cabling for Parkend. I gave up working at Parkend when I found the electricity not available upstairs for a mains soldering iron. I went home to work on the second relay set for Parkend, returning with Jean to have lunch with Rick and Peter at the Cycle shop cafe.
25 July : Rick and Martin went off looking for problems on the Parkend junction pairs. They found some of course but as usual did not manage to get the two lost junctions back into use. Peter, Sam and Alan ran in yet more data cabling at Parkend signal box. I changed a capacitor in the multi level relay set at Parkend so that the output pulse speed was correct. Then I got on with wiring the head on the concentrator. A slow job, finding the wires, terminating them and then testing the result. I got about halfway through the job before we packed in for the day.
27 July : Rick came in today to chase Norchard to Parkend junctions. He is documenting the situation as he goes as the position appears to change every time we try to fault find on the Norchard to Whitecroft section. The position at the end of the day was not encouraging. Junctions 1 and 2 outgoing from Norchard are both dis one leg between DPs 3X and 3Y. These two junctions proved to be OK at the last investigation so something strange is happening. Junction 3 seems OK between Norchard and DP3Z but for now is still busied out and needs checking on Wednesday. Junction 4 had a broken wire in DP3Z. It is repaired and put back into service. Subsequently I dug out the testers 301 that we have, checked them and delivered them to Rick so that he can return to start locating the DP3X to DP3Y fault.
28 July : This morning, Peter and Roger came in at ten. Roger worked on programming the ATAs at Parkend. Peter mounted woodwork, new trunking and the fuse cabinet. Sam met them at Norchard for lunch, following which Sam terminated the cat5 upstairs, whilst Peter sorted out the patch panel end. Peter and Sam then spent time cleaning and tidying our area of the box. S&T have cleared away their shelving by the door, which gives us much more space!
1 August : Sam was in for the morning attempting to tidy our office, it needs it. Rick went lineside tracking intermittent disconnections on the cable to Whitecroft. He ended the morning with all four junctions back in service, but for how long? Charles and Alan were trying to fit a data cabinet to the wall in the Parkend staff room. I fitted the new multi level relay set in Parkend exchange and tested it. Just one fault, a wire I had omitted during construction. After that the set seems to run very well. Peter and I could not test the relay set to a digital connection as the broadband was not working but I managed to check it over a strowger connection. In the evening the group, with partners, went on the fish and chip DMU trip. Very good fish and chips from the Bream chippy.
4 August : Roger and I spent the morning attempting to get the strowger exchange to send its digits correctly to the digital Asterisk system. We ended with the digits entering the Asterisk correctly but with the Asterisk not knowing what to do with them. The diary needs an update on this. Sam and Peter have just about got the digital circuits in the station building working back to the new data cabinet in the staff room. Perhaps we are close to moving elsewhere to do some telecoms work. We went to lunch at Kaplan's.
7 August : Rick and Martin got the replacement DP3X mounted and the cables changed over. It still has to be tidied and completed.
8 August : The token was out of action so Rick and Martin continued to work on the Norchard to Whitecroft cable problems. They at least got the token back working by swapping out the pair. Peter is surveying Lydney Junction for a new MDF for use with the S&T changes in conjunction with the crossing renewals. I continued on with the Parkend concentrator wiring. It's nearly finished. 11 August : Roger and Peter worked at Parkend on the Asterisk system. The Raspberry Pi was updated with a new IP address so it was not dependent on the DHCP server and 4710 brought into service (digital phone in the Signal Box). Further testing was carried out with John's relay set and re-generator that links Strowger to a Grandstream HT802 ATA and then into the Raspberry Pi. The ATA receives the pulse train correctly 90% of the time but occasionally calls to 402 are processed as calls for 492 so there is an issue still to resolve with Loop Disconnect dialling.
Roger added ext4399 and ext4799 to the Pi which generates the UK Inverted Ring sound, transferred from Paul's config at Norchard. Peter tested the ATAs from the Krone blocks and updated the paper
records for Parkend. We also found one junction from Asterisk in Norchard to Strowger in Norchard was down (every 2nd call completed). One to look at next week. 15 August : A good day for the Telecoms Group. Rick and Martin finally tracked down the last fault on the Norchard to Whitecroft cables and have got all the junctions in service. Peter, Charles and Alan made an excellent start on fitting the new Lydney Junction MDF and the associated cabling. I finished the wiring of the Parkend concentrator and as far as I can see it now works. It needs two to test it properly next week and it needs a new armature on one relay which is holding in after being operated. Nice for everyone to go home feeling that something positive has been accomplished.
8 August : The token was out of action so Rick and Martin continued to work on the Norchard to Whitecroft cable problems. They at least got the token back working by swapping out the pair. Peter is surveying Lydney Junction for a new MDF for use with the S&T changes in conjunction with the crossing renewals. I continued on with the Parkend concentrator wiring. It's nearly finished.
11 August : Roger and Peter worked at Parkend on the Asterisk system. The Raspberry Pi was updated with a new IP address so it was not dependent on the DHCP server and 4710 brought into service (digital phone in the Signal Box). Further testing was carried out with John's relay set and re-generator that links Strowger to a Grandstream HT802 ATA and then into the Raspberry Pi. The ATA receives the pulse train correctly 90% of the time but occasionally calls to 402 are processed as calls for 492 so there is an issue still to resolve with Loop Disconnect dialling. Roger added ext4399 and ext4799 to the Pi which generates the UK Inverted Ring sound, transferred from Paul's config at Norchard. Peter tested the ATAs from the Krone blocks and updated the paper records for Parkend. We also found one junction from Asterisk in Norchard to Strowger in Norchard was down (every 2nd call completed). One to look at next week.
15 August : A good day for the Telecoms Group. Rick and Martin finally tracked down the last fault on the Norchard to Whitecroft cables and have got all the junctions in service. Peter, Charles and Alan made an excellent start on fitting the new Lydney Junction MDF and the associated cabling. I finished the wiring of the Parkend concentrator and as far as I can see it now works. It needs two to test it properly next week and it needs a new armature on one relay which is holding in after being operated. Nice for everyone to go home feeling that something positive has been accomplished.
Current Value of our Fund
Return to the Diary Menu